# MoDA: Multi-modal Diffusion Architecture for Talking Head Generation

### v1

@@ 4.2.2 Qualitative Evaluation
As shown in Fig. 5, we selected two types of cases from the wild dataset for visual comparison. For each character, we generated videos using each method and selected frames from the same location for comparison. Analyzing the results, previous one-stage-based methods, including Hallo2, Hallo, and JoyHallo, suffer from appearance blurring and unnatural expressions during temporal inference due to the strong entanglement between appearance and motion. In contrast, our proposed two-stage framework effectively mitigates these issues, ensuring high consistency in the generated details. Compared to the two-stage method JoyVASA, MoDA generates richer expressions, better lip synchronization, and more natural head movements, thanks to our multi-modal motion generation network, which effectively integrates deep information across different modalities.
@@
Method FVD $\downarrow$ Sync-C $\uparrow$ Sync-D $\downarrow$
w/o C2F 221.631 5.535 8.465
w CABA 232.291 5.331 8.692
Full Model 214.182 5.757 8.346
4.2.3 Visualization Results in Complex Scenarios.
We further investigate MoDA’s generation performance in complex scenarios. Specifically, for the visual modality, we utilize portraits of both real humans and animated characters, each paired randomly with various audio types, including speech, singing, recitation, and others. As shown in Fig. 6, our method demonstrates strong performance in various complex scenarios. In addition, it is capable of efficient long-duration inference and fine-grained facial expression control.
@@ 4.3 Ablation Study
To validate the effectiveness of our method components, we deactivate each of them and report result in Table 3. For more ablation experiments and implementation details, please refer to the supplementary materials.
@@ 4.3.1 Cross-attention
We replace our model with a Cross-Attention-Based Architecture (CABA) to perform multi-modal fusion for evaluation purposes. As shown in Table 3, this substitution results in a performance drop in all metrics. Moreover, as illustrated in Fig. 7, the subject exhibits mouth-closing failures when lacking deep cross-modal interaction. To investigate this, we conducted two experiments under the same settings used in the w/ CABA variant as shown Fig. 7: 1) replacing the audio with the same image did not resolve the mouth closure issue (Replace audio); 2) replacing the image with the same audio addressed the issue (Replace image). These results suggest that cross-attention fails to adapt audio features to different identity conditions. In contrast, MoDA enables dynamic adaptation of audio features based on contextual factors such as motion, identity, and emotion.
4.3.2 Coarse-to-Fined Feature Fusion and Interaction
We replace the Coarse-to-Fine Feature Fusion with a fully four-stream architecture (w/o C2F). As shown in Table 3, this design leads to performance degradation, particularly in synchronization and motion quality, emphasizing the necessity of progressive feature fusion for effective cross-modal integration. These results indicate that independently learning weights introduces substantial redundant parameters, resulting in convergence challenges (with 904M parameters in the w/o C2F variant). In contrast, by gradually sharing weights, the proposed coarse-to-fine strategy significantly reduces the parameter count to 370M, leading to both greater efficiency and improved performance.
@@ 4.4 Discussion
A key limitation of our framework lies in the performance constraints introduced by the first-stage model, including the appearance encoder, motion encoder, and decoder. In particular, the Liveportrait model used at this stage performs poorly when handling large pose variations, leading to a degradation in output quality. Moreover, it often produces unnatural distortions when processing subjects with complex head accessories. These issues may arise from limitations in the training data or intrinsic constraints of the GAN-based generator architecture.
To address these issues, future work will explore more robust first-stage generation techniques. We plan to improve the training dataset by incorporating a wider range of head pose variations and complex head accessories. In addition, we will investigate more advanced video generation models [45, 18, 15, 22] to further improve the quality and expressiveness of the result.
@@ 5 Conclusion
We present MoDA, a two-stage multimodal diffusion architecture specifically designed for one-shot talking-head synthesis. Our method effectively leverages multimodal information to map audio signals to motion sequences in an identity-agnostic latent space, followed by a dedicated face generator to produce high-quality videos. MoDA can synthesize natural and coherent talking head videos from a single portrait image with high efficiency, while also supporting fine-grained expression control. It significantly outperforms previous state-of-the-art methods in terms of video quality, motion diversity, and naturalness.

### v2

@@ 4.2.2 Qualitative Evaluation
As shown in Fig. 5, we selected two types of cases from the wild dataset for visual comparison. For each character, we generated videos using each method and selected frames from the same location for comparison. Analyzing the results, previous one-stage-based methods, including Hallo2, Hallo, and JoyHallo, suffer from appearance blurring and unnatural expressions during temporal inference due to the strong entanglement between appearance and motion. In contrast, our proposed two-stage framework effectively mitigates these issues, ensuring high consistency in the generated details. Compared to the two-stage method JoyVASA, MoDA generates richer expressions, better lip-synchronization, and more natural head movements, thanks to our multi-modal motion generation network, which effectively integrates deep information across different modalities.
4.2.3 Visualization Results in Complex Scenarios.
We further investigate MoDA’s generation performance in complex scenarios. Specifically, for the visual modality, we utilize portraits of both real humans and animated characters, each paired randomly with various audio types, including speech, singing, recitation, and others.
@@
Method FID $\downarrow$ FVD $\downarrow$ Sync-C $\uparrow$ Sync-D $\downarrow$
w/ CABA 47.365 232.291 5.331 8.692
w/o C2F 44.548 221.631 5.535 8.465
w/ MAF 48.358 216.982 5.511 8.527
Full Model 44.071 205.442 5.878 8.135
As shown in Fig. 6, our method demonstrates strong performance in various complex scenarios. In addition, it is capable of efficient long-duration inference and fine-grained facial expression control.
@@ 4.3 Ablation Study
For more ablation experiments and implementation details, please refer to the supplementary materials.
@@ 4.3.1 Cross-attention
We replace the MoDA with a Cross-Attention-Based Architecture (CABA) to perform multi-modal fusion for evaluation purposes. As shown in Table 3, this substitution results in a performance drop in all metrics. Moreover, as illustrated in Fig. 7, the subject exhibits mouth-closing failures when lacking deep multi-modal interaction. To investigate this, we conducted two experiments under the same settings used in the w/ CABA variant as show in Fig. 7: 1) replacing the audio with the same image did not resolve the mouth closure issue (Replace audio); 2) replacing the image with the same audio addressed the issue (Replace image). These results suggest that cross-attention fails to adapt audio features to different identity conditions. In contrast, MoDA enables dynamic adaptation of audio features based on contextual factors such as motion, identity, and emotion.
4.3.2 Coarse-to-Fined Feature Fusion
We replace the Coarse-to-Fine Feature Fusion with a fully four-stream architecture (w/o C2F). As shown in Table 3, this design leads to performance degradation, particularly in synchronization and motion quality, emphasizing the necessity of progressive feature fusion for effective cross-modal integration. These results indicate that independent learning of weights introduces substantial redundant parameters, resulting in convergence challenges (with 904M parameters in the w/o C2F variant). In contrast, by gradually sharing weights, the proposed coarse-to-fine strategy significantly reduces the parameter count to 370M, leading to both greater efficiency and improved performance.
To further investigate the effectiveness of the fusion strategy, we modified the original dual-stream design by creating an MAF variant that first merges motion with the auxiliary cues (emotion + identity) into one branch, while audio stays in the other. Table 3 also shows that MAF performs worse than even the w/o C2F baseline. Directly fusing the heterogeneous noisy motion and auxiliary condition features confuses the network. Our full model first merges audio with the auxiliary cues; speech audio inherently carries emotion and identity signals, providing a natural bridge and delivering better accuracy and efficiency.
@@ 4.4 Discussion
For the disentanglement stage, we can use alternatives such as MegaPortraits or Face Vid2Vid, but like LivePortrait they degrade on large pose changes or when the subject wears complex head accessories. This may be due to limitations in the training data or intrinsic constraints of the GAN-based architecture. To address these challenges, we plan to improve the training dataset by incorporating a wider range of variations in head poses and complex head accessories. Furthermore, we will investigate more advanced video generation models [16, 20] to improve the quality and expressiveness of the generated results.
@@ 5 Conclusion
We propose MoDA, a two-stage multi-modal diffusion framework for one-shot talking head generation. MoDA effectively leverages multi-modal information to map audio to motion sequences in an identity-agnostic latent space, which are then translated into video frames by a pre-trained face renderer. From a single portrait, MoDA produces high-quality, expressive, and controllable talking head videos, surpassing previous methods in quality, diversity, and naturalness with high efficiency.